The average cholesterol content of a certain can goods is 215 milligrams and the cholesterol deviation is 15 milligrams.assume the variable is normally distributed. If a sample of 25 canned goods is selected, what is the probability that the mean of the sample will be greater than 220 miligrams?


1
Expert's answer
2021-06-24T18:35:25-0400

Let "X=" the mean of the sample: "X\sim N(\mu, \sigma^2/n)."

Given "\mu=215\ mg, \sigma=15\ mg, n=25"


"P(X>220)=1-P(X\leq 220)"

"=1-P(Z\leq \dfrac{220-215}{15/\sqrt{25}})\approx1-P(Z\leq 1.6667)"

"\approx0.04779"